When calcium carbonate is added to hydrochloric acid, calcium chloride, carbon dioxide, and water are produced.

CaCO3(s) + 2HCl(aq) ⟶ CaCl2(aq) + H2O(l) + CO2(g)

1. How many grams of calcium chloride will be produced when 29.0 g of calcium carbonate is combined with 15.0 g of hydrochloric acid?

2. Which reactant is in excess? CaCO3 or HCl

3. How many grams of the excess reactant will remain after the reaction is complete?
1
Expert's answer
2019-10-07T08:31:50-0400

n(CaCO3) = 29.0 / 100.1 = 0.29 mol

n(HCl) = 14/ 36.5 = 0.384 mol

m(CaCl2) = 111.1 * 0.384 / 2 = 21.3 g

m(CaCO3) = (0.29 – 0.384 / 2) * 100.1 = 9.8 g

Answer:

1) 21.3

2) Calcium carbonate is in excess.

3) 9.8 g grams of calcium carbonate